OpenStudy (anonymous):

If a pro basketball player has a vertical leap of about 35 inches, what is his hang time? Use the hang-time function V=48T^2

OpenStudy (anonymous):

what does V stand for in the formula?

OpenStudy (anonymous):

I think vertical? Not sure..this week in class I have been very confused....

OpenStudy (anonymous):

if it's the vertical distance, then substitute it , 35=48T^2.

OpenStudy (anonymous):

\[35=48T^2\]\[T=\sqrt{35/48}\] T=.8539 seconds

OpenStudy (anonymous):

yep, that may be right. but what if V turns out to be VELOCITY, though not so likely?

OpenStudy (anonymous):

It is for vertical leaps

OpenStudy (anonymous):

then .8539 is the answer

OpenStudy (anonymous):

then .8539 is the answer

OpenStudy (anonymous):

I just tried it and it kicked back saying it was wrong as well...cause my answer was close to that as well

OpenStudy (anonymous):

the V means vertical leaps in minutes

OpenStudy (anonymous):

I mean inches

OpenStudy (anonymous):

Please help me...

OpenStudy (anonymous):

\[\sqrt{9x+67}=x+5\]

OpenStudy (anonymous):

I need this solved...I really am just checking the answers I got wrong to see where I went wrong

OpenStudy (anonymous):

x = 6